Determine the elements of order 15 of U(225)


1
Expert's answer
2021-09-21T02:54:39-0400

We can write "U_{225}\backsimeq U_{15}\times U_{15}" [By chinese remainder theorem]

Hence any element of "U_{15}"  has orders 1,2 or 4. Since  "|U_{15}|=\u03c6(15)=8"

there is no element of "U_{15}"  that generates it, hence "U_{15}"  is not a cyclic group.

So, we can write:

"U(225)=U(15^2 )=U(15)=1,2,4,7,8,11,13,14.\\
The \ cyclic \ subgroups \ generated \ by \ these \ elements \ are:\\
<2>=<8>=1,2,4,8,\\
<4>=1,4,\\
<7>=1,4,7,13,\\
<11>=1,11(note \ that \ 11=−4),\\
<14>=1,14(note \ that \ 14=−1),\\
<1>=1.\\
Observe \ that \ U(15) \ cannot \ be \ cyclic \ since \ it \ has \ three \\ subgroups \ of \ order \ 2 \ and \ two \ of \ order \ 4."